The nanda 11 500 used Aw 229 009 28
Answer #1
Answer to Part (a)
Current Assets / Current Liabilities = Current Ratio
2016 184000 / 130000 = 1.42
2015 216000 / 93000 = 2.32
Answer to Part (b)
Cash + Short Term Investments + Net Receivables / Current Liabilities = Acid Test Ratio
2016 26000 + 8000 + 50000 / 130000 = 0.65
2015 32000 + 29000 + 73000 / 93000 = 1.44